# How Good Are We? Evaluating Cell AI Foundation Models in Kidney Pathology with Human-in-the-Loop Enrichment

### v1

@@ 5.2 Data-Enriched Foundation Models Fine-Tuning
To enhance model performance while minimizing reliance on pixel-level human annotation, we further fine-tuned the foundation models using our data enrichment strategies, which ensemble the evaluation outcomes of multiple foundation models. For each foundation model, the pretrained weights served as the baseline method. First, we evaluated the baseline performance of each cell foundation model using our hold-out testing dataset. The results of assessing the baselines for Cellpose, StarDist, and CellViT are shown in blue in Table. 2. Consistent with our previous performance evaluation from human feedback, CellViT demonstrates superior performance compared to the other two models, with an F1 score of 0.7838, followed by StarDist at 0.7380, and Cellpose at 0.6748. Although StarDist achieves a high precision score of 0.9158, its lower recall of 0.6331 suggests that it misses certain amount of kidney cell nuclei. Cellpose exhibits the largest domain gap with the lowest F1 score and shows at least 10% decrease in all evaluation metrics (F1 score, Precision, and Recall) compared to CellViT.
Then, we investigated the performance of foundation models from our data-enriched fine-tuning strategies. In Fig. 5, We grouped the fine-tuning results across all three foundation models in this work. Each column represents the fine-tuning results for an individual foundation model, while each row compares an instance segmentation performance metric (F1 score, Precision, or Recall) across the models. Different colors represent distinct fine-tuning strategies or settings. The X-axis in each plot corresponds to the incremental percentage (25%, 50%, 75%, and 100%) of the training dataset used for each fine-tuning strategy, and the Y-axis shows the performance metrics. From the fine-tuning performance trends illustrated in Fig. 5, both the Cellpose and StarDist models demonstrate significant improvements in F1 score compared to their baseline models. These gains highlight the effectiveness of the data enrichment strategies employed during fine-tuning. The details of instance segmentation performance gains are provided below with the evaluation metric values referenced in Table. 2.
  1. 1.
    For the Cellpose model, fine-tuning with only “good” image patches at all training dataset scales significantly boosts performance compared to the model baseline, with the F1 score increasing from 0.6748 to 0.7529, Precision improving from 0.7006 to 0.8447, and Recall increasing from 0.6694 to 0.6986. However, fine-tuning with “bad” image patches alone results in lower performance. Although Precision reaches high at 0.8401, Recall drops significantly, with the highest recall being 0.5592, which is over 10% below the baseline recall score. The F1 scores also decrease, with the highest value being 0.6444 below baseline F1 score at 0.6748, indicating poor performance when relying solely on these “hard” samples. Fine-tuning with a combination of “good” and “bad” image patches yields stable improvements across all metrics. This shows that the presence of “good” image patches from multiple foundation models’ predictions is crucial, which compensates for the model’s inability to learn effectively from “bad” image patches alone. While the F1 scores and Precision are slightly lower compared to using only “good” patches, they still significantly outperform the baseline.
  2. 2.
    For the StarDist model, fine-tuning trend in Fig. 5 across all data enrichment fine-tuning strategies resulted in notable instance segmentation performance gains. The F1 score increased significantly from the baseline of 0.738 to 0.8229, and Recall improved from 0.6331 to all above 0.7802. While Precision saw a slight drop from 0.9158 to around 0.89, the decrease is minimal compared to the substantial improvements in F1 and Recall. This highlights the effectiveness of each data enrichment strategy in improving segmentation quality, with “bad” samples minimizing the knowledge gap through expert labeling, and “good” samples distilling knowledge from other foundation models. This demonstrates that our HITL design, based on the curation outcomes of multiple foundation models, can effectively improve model performance.
Lastly, according to both Fig. 5 and Table. 2, CellViT shows relatively consistent F1 scores across all fine-tuning strategies. Fine-tuning with only “good” data slightly increases Precision but causes a small reduction in Recall (e.g., from 0.7629 to 0.7476 with 100% “good” data, with the highest recall being 0.7621). In contrast, fine-tuning with only “bad” data improves Recall (from 0.7629 to 0.8137) but reduces Precision (from 0.8233 to 0.7898). Combining “good” and “bad” data offers the best of both data enrichment strategies, with improved F1 scores at every training scale. For instance, using 100% “good” and “bad” image patches achieves the highest F1 score (0.7952), with balanced Precision (0.8302) and Recall (0.7731), all surpassing the baseline. Although the performance gain is not as pronounced as with Cellpose or StarDist, integrating both “good” and “bad” image patches results in the most balanced overall performance, enhancing both Precision and Recall.
Summarizing the best fine-tuned model for each cell foundation model, Fig. 6 compares these models to their corresponding baselines. This indicates that all models enhance segmentation performance after fine-tuning. Notably, the baseline model with the highest F1 score before fine-tuning (CellViT) does not yield the best segmentation outcomes after fine-tuning. In contrast, StarDist benefits the most from the proposed data enrichment strategies, achieving the highest F1 score of 0.8229.
Figure 7: Impact of Incrementally Increasing Percentage (%) of High-Quality Labels on Model Performance. The dashed line represents the baseline comparison setting, using 25% pseudo labels from multiple foundation models’ “good” predictions and 25% pathologist-corrected labels, which represent the gap of the models.
@@ 5.3 Effect of Label Imbalance on Model Performance
Following the quantitative analysis above, compared to Cellpose and StarDist, which show a significant increase in performance over their baseline methods, CellViT’s overall performance gain is not as significant. We hypothesized that this can be attributed to label imbalances in the training dataset. As shown in Fig. 3, CellViT generates the majority of the pseudo labels, leading to an imbalance between these pseudo labels and pathologist-corrected labels (representing “bad” image patches), which highlights the domain gap. Additionally, there is an imbalance within the pseudo labels themselves; the proportion of pseudo labels rated as “good” by other models, while not by CellViT, is less pronounced compared to the fine-tuning of StarDist and Cellpose. This imbalance can be observed by examining the difference in the number of “good” image patches between the fused model and the individual foundation models in Fig. 3. A larger difference indicates that data enrichment distills more image patches needed by the model.
Thus, we performed an ablation study to assess the impact of adding high-quality labels, which the foundation models need in our imbalanced fine-tuning dataset. For ease of implementation, we incrementally increased the percentage of pathologist-corrected “bad” image patches in the training dataset. The baseline comparison (dash line) used 25% pseudo labels from multiple models’ “good” predictions and 25% pathologist-corrected labels. In this study, we incrementally incorporated an additional 25% of “bad” image patches into the training dataset, presenting the model performance results as blue curves in Fig. 7. Consistent with our previous analysis, we observe that the inclusion of “bad” image patches, which are often faintly stained samples, decreases the performance of Cellpose. This suggests that fine-tuning the Cellpose model requires “good” image patches to maintain the pretrained knowledge and avoid catastrophic forgetting during our continual fine-tuning. For both CellViT and StarDist, we observe a clear pattern of increased performance over the baseline as the availability of model-required high-quality labels rises. Incorporating additional “bad” image patches leads to notable increases in the F1 score. For CellViT, when the percentage of “bad” image patches in the training dataset is increased to 50%, CellViT achieves an F1 score of 0.795, which is equivalent to its peak performance when using all “good” and “bad” image patches, as shown in Table 2. With inclusion of all “bad” image patches, it reaches highest F1 score of 0.798. This improvement is attributed not only to the inclusion of high-quality labels but also to the oversampling of these image patches in the training dataset. Overall, the results demonstrate that the data enrichment strategy of utilizing both “good” and “bad” image patches together enhances the performance of StarDist and CellViT, and oversampling the model-needed data is a crucial aspect of our approach.
@@ 5.4 Visualization of Improved Model Performance
In this section, we present the qualitative visualization results of our enhanced kidney cell nuclei instance segmentation, evaluated using the best fine-tuning strategies from three foundation models. Three example image patches (A, B, and C) are presented in Fig. 8. The upper panel for each example shows the baseline performance across all models, while the lower panel displays the segmentation performance following data-enriched fine-tuning. As shown in Fig. 8(A), our fine-tuned Cellpose model improves the instance segmentation on long and flat nuclei. (B) highlights the improvement of our StarDist model in segmenting dense nuclei in a glomerulus. The last panel shows the improvement in segmenting nuclei in PAS-stained images, which are underrepresented in the current histology dataset. Most faintly stained nuclei that were missing in the original predictions within a glomerulus are detected by the fine-tuned StarDist and CellViT models. In contrast, the fine-tuned Cellpose model generates fewer false positive predictions compared to its baseline. Additionally, as illustrated in both (A) and (B), the highlighted regions in the image patches represent knowledge gaps in the target domain where the inferior model struggles with certain types of image patches, while the other models do not. Examples of our fine-tuning results demonstrate that this less effective model can benefit from the knowledge of the other models to improve its performance. Additional visualizations of the qualitative results (D)-(L) are provided in Supplementary Figs. S1, S2, and S3 (See Supplementary Information for additional data)
@@ 6 Discussion
The evaluation of Cellpose, StarDist, and CellViT reveals that a performance gap still exists between general and kidney-specific nuclei segmentation. Cell nuclei segmentation cannot be considered a fully solved problem across all organs, tissues, and imaging modalities when utilizing off-the-shelf foundation models. Our foundation model-based data enrichment strategies aim to meet the need for a kidney-level cell foundation model and to reduce labeling costs. In this process, the collective image curation from multiple foundation models is crucial, as a single model may not capture the diversity of segmentation styles in the target kidney domain. By leveraging the versatility of multiple foundation models, annotation efforts are minimized and transformed into time-efficient image rating curation, which can be further enhanced through the model’s uncertainty analysis. In the first round of HITL fine-tuning, all three foundation models demonstrated improved performance over their baselines, with StarDist achieving the highest segmentation performance, achieving an F1 score of 0.82. Both StarDist and CellViT highlight that including “good” and “bad” image patches is the most effective training strategy in our approach. Although Cellpose shows a significant increase in performance (with the F1 score rising from 0.67 to 0.75) with our fine-tuning, it demonstrates lower efficacy when only hard image samples are included. This is primarily because the Cellpose model is designed for fluorescently labeled images, whereas the other two models have been adapted for H&E-stained images. For the training of segmentation task, 512 $\times$ 512 images were converted into grayscale nuclear bands, which can lead to information loss, particularly for hard images with very light staining colors and complex tissue morphological patterns.
A limitation of this approach is that it requires more expert knowledge and image preprocessing to enhance nuclear intensity while minimizing the noisy background in kidney whole-slide images. Another aspect that requires further improvement is the imbalance of labels in the training dataset. As investigated in our ablation study, oversampling the labels that highlight the domain gap is essential. However, this approach presents a trade-off, as labeling or curating these samples from a large-scale dataset incurs additional manual costs. In our future research, we plan to employ semi- or fully-automated model selection to help mitigate this workload.
@@
Lastly, as the models become better adapted to the specific kidney organ and the performance gap decreases among three models after the first round of fine-tuning, we anticipate generating improved curation results, including a higher number of “good” ratings, a reduction in “bad” samples, and more accurate prediction masks in the next round of HITL fine-tuning. This improvement can be further enhanced by utilizing our improved model (e.g., StarDist) in QuPath software for label correction by pathologists.
7 Conclusion
In this work, our evaluation on a comprehensive kidney dataset indicates that cell nuclei segmentation in histopathology still needs further improvement with more specific organ-targeted foundation models. Among the evaluated models, CellViT demonstrated superior performance in segmenting nuclei in kidney pathology, achieving an F1 score of 0.78 in our hold-out testing dataset. To enhance the performance of foundation models for segmenting kidney nuclei, we utilized multiple foundation models in a collective rating-based curation process and used the results as data enrichment strategies during fine-tuning. Our experimental results show that all three foundation models improved over their baselines, with StarDist achieving the highest F1 score of 0.82. Both StarDist and CellViT demonstrate that using “good” and “bad” image patches is the most effective training strategy. Notably, Cellpose, primarily a fluorescence-based model, raised its F1 score from 0.67 to 0.75 by incorporating “good” predictions from the other models. This mutual enhancement highlights the effectiveness of our strategy in improving segmentation outcomes. The improved model can further be applied in QuPath software to facilitate more efficient workflows in clinical kidney pathology.
Disclosures
The authors of the paper have no conflicts of interest to report.
Code, Data, and Materials Availability
@@ 7 Conclusion
Code will be coming soon and publicly available at https://github.com/hrlblab/AFM_kidney_cells.
Data will be provided after satisfying Vanderbilt University and Vanderbilt University Medical Center’s data use agreement.
@@ Acknowledgments
This work was supported by the National Institutes of Health under award numbers R01EB017230, R01DK135597, T32EB001628, K01AG073584, and 5T32GM007347, and in part by the National Center for Research Resources and Grant UL1 RR024975-01. This study was also supported by the National Science Foundation (1452485, 1660816, and 1750213). The Vanderbilt Institute for Clinical and Translational Research (VICTR) is funded by the National Center for Advancing Translational Sciences (NCATS) Clinical Translational Science Award (CTSA) Program, Award Number 5UL1TR002243- 03. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH or NSF. This work was also supported by Vanderbilt Seed Success Grant, Vanderbilt Discovery Grant, and VISE Seed Grant. We extend gratitude to NVIDIA for their support by means of the NVIDIA hardware grant. This work was also supported by NSF NAIRR Pilot Award NAIRR240055.
@@
Supplementary Information
Supplementary Figures S1, S2, and S3. Additional Qualitative Results of Enhanced Kidney Cell Nuclei Instance Segmentation. In addition to (A)-(C) shown in the Fig. 8, (D)-(L) are provided in Supplementary Figures S1 to S3.
Biographies and photographs of the other authors are not available.

### v2

@@ 3.2 Data-Enriched Foundation Models Fine-Tuning
Baseline Performance. First, We used the pretrained weights of each foundation model as the baseline and evaluated their performance on our hold-out test set; results are shown in blue in Table 2. Consistent with prior human feedback, CellViT outperforms the other two models, with an F1 score of 0.7838, followed by StarDist at 0.7380, and Cellpose at 0.6748. Although StarDist achieves a high precision score of 0.9158, its lower recall of 0.6331 suggests that it misses certain amount of kidney cell nuclei. Cellpose exhibits the largest domain gap with the lowest F1 score and shows at least 10% decrease in all evaluation metrics compared to CellViT.
Fine-tuned Performance. Then, we investigated the performance of foundation models from our data-enriched fine-tuning strategies. The details of instance segmentation performance gains are provided below with the evaluation metric values referenced in Table 2.
(1) Cellpose: As shown, training on “hard” patches alone significantly degrades performance, particularly in the F1 score and Recall. In contrast, the inclusion of large-scale “easy” (pseudo-labeled) training samples consistently improves performance, with optimal results achieved using only “easy” patches, significantly boosting all metrics over the baseline (F1: 0.6748 $\rightarrow$ 0.7529).
(2) StarDist: All fine-tuning strategies lead to strong instance segmentation performance gains, especially in F1 (0.738 $\rightarrow$ 0.8229) and Recall (0.6331 $\rightarrow$ 0.7802). While Precision saw a slight drop from 0.9158 to around 0.89, the decrease is minimal compared to the substantial improvements in F1 and Recall. “Easy” patches support knowledge distillation, while “hard” ones close domain gaps, confirming the effectiveness of HITL-guided data enrichment.
(3) CellViT: Performance remains stable across strategies. Combining both “easy” and “hard” patches yields the highest F1 score (0.7952) and the most balanced results, though the gains are smaller than those observed for Cellpose or StarDist. This may be due to noise introduced by pseudo-labels from less accurate models, which will be examined in the next section 3.3.
Optimal Annotation Strategies. Lastly, we summarize the optimal training and annotation strategies for the three foundation models in Fig. 6. We use the F1-score as the primary evaluation metric, as it balances both precision and recall. Among the models, the Fluorescent-dominant model, Cellpose, achieves optimal performance when trained with only “easy” labels. Its performance degrades with the inclusion of “hard” labels, likely due to their significant domain difference from the model’s pre-training data. In this context, “hard” cases are the training cases where automatic segmentation fails human QA and employ pixel-level manual corrections, while “easy” cases are those where the automatic results successfully pass human QA without modification. In contrast, the H&E-dominant models consistently benefit from incorporating human-corrected labels (“hard”), outperforming their respective baselines. Specifically, StarDist achieves the best performance with “hard” labels (attaining the highest F1 score of 0.8229), while CellViT performs optimally when trained with both “easy” and “hard” labels. In summary, the combined use of “easy” and “hard” patches proved to be the most effective and broadly applicable annotation strategy, consistently enhancing performance across all three models.
Choice of training dataset size. In this study, we selected common proportions—25%, 50%, 75%, and 100%—to evaluate whether the impact of annotation strategies across the three models remains consistent under different dataset sizes. Given that the annotation efforts are limited to only a total of 198 “hard” labels and a majority comes from pseudo-labels. We hypothesize that maintaining the overall scale of the training dataset to include enough representative samples is important. To further assess this, we conducted additional experiments using only 10% of “easy” labels (N = 1,181), “hard” labels (N = 20), and their combination to validate the results. The outcomes are also presented in Table 2.
As shown in Table 2, reducing the training dataset to 10% results in smaller performance gains across all models, though the overall trend remains consistent—particularly for the CNN-based models (Cellpose and StarDist), as illustrated in Fig. 6. In contrast, performance drops more noticeably below the baseline for the ViT-based CellViT. With only 20 “hard” samples, fine-tuning becomes challenging, making comparisons between groups (e.g., “easy” vs. “easy” + “hard”) less reliable. To broadly improve model performance, a sufficient amount of both “hard” and “easy” data (e.g., at least 25%) is essential for effective fine-tuning. Alternatively, maintaining segmentation accuracy with smaller datasets requires additional data curation to ensure representative coverage.
Robustness. To better evaluate the robustness of the experiments, we applied 5-fold cross-validation and compared F1-scores as part of an ablation study. As shown in Supplementary Table S2, the results compare model performance under different training annotation strategies.
@@ 3.3 Effect of Label Imbalance on Model Performance
Compared to Cellpose and StarDist, CellViT shows a smaller performance gain over its baseline, which we hypothesize is due to imbalanced annotation quality. (1) variation in pseudo-label quality introduces noise. As shown in Fig. 4, CellViT—the strongest baseline model—generates the majority of pseudo labels; however, incorporating pseudo labels from weaker models may introduce noise that hinders its fine-tuning. (2) There is an imbalance between pseudo labels (from “easy” cases) and gold-standard annotations (from “hard” cases). To investigate these effects, we conducted an ablation study in which we incrementally added noise-free, pathologist-corrected “hard” image patches to the training set, as shown in Fig. 7.
Consistent with Fig. 6, the Fluorescent-dominant Cellpose model shows decreased performance when “hard” patches are included, although the annotations are gold-standard, but the usually faintly stained patches are too different from its pre-training dataset. In contrast, the H&E-dominant models, CellViT and StarDist, show consistent performance improvement as more high-quality “hard” annotations are added. The differing trend may be attributed to the training data used for each model: the Cellpose model was predominantly trained on fluorescence microscopy images with Fluorescent staining (a “Fluorescent-dominant” model), while CellViT and StarDist were primarily trained on histology images with H&E staining (forming “H&E-dominant” models). Since the “hard” dataset is composed mostly of H&E-stained images, this domain gap likely contributes to the observed differences in performance. Notably, CellViT achieves its highest F1 score of 0.798 when all “hard” image patches are included. This aligns with the optimal annotation strategies shown in Fig. 6, where the combined use of “easy” (pseudo-labeled) and “hard” (pathologist-corrected) patches proves to be the most effective and broadly applicable approach, consistently improving performance across all three models.
Figure 7: Impact of Incrementally Increasing Percentage (%) of Noise-Free, Pathologist-Corrected “Hard” Labeled Data on Model Performance. As shown, the inclusion of “hard” labeled images decreases the performance of Cellpose. In contrast, for both CellViT and StarDist, incorporating these data leads to increased F1 scores.
@@ 3.4 Visualization of Improved Model Performance
In this section, we present the qualitative visualization results of our enhanced kidney cell nuclei instance segmentation, evaluated using the best fine-tuning strategies from three foundation models. Three example image patches (A, B, and C) are presented in Fig. 8. The upper panel for each example shows the baseline performance across all models, while the lower panel displays the segmentation performance following data-enriched fine-tuning. As shown in Fig. 8(A), our fine-tuned Cellpose model improves the instance segmentation on long and flat nuclei. (B) highlights the improvement of our StarDist model in segmenting dense nuclei in a glomerulus. The last panel shows the improvement in segmenting nuclei in PAS-stained images, which are underrepresented in the current histology dataset. Most faintly stained nuclei that were missing in the original predictions within a glomerulus are detected by the fine-tuned StarDist and CellViT models. In contrast, the fine-tuned Cellpose model generates fewer false positive predictions compared to its baseline. Additionally, as illustrated in both (A) and (B), the highlighted regions in the image patches represent knowledge gaps in the target domain where the inferior model struggles with certain types of image patches, while the other models do not. Examples of our fine-tuning results demonstrate that this less effective model can benefit from the knowledge of the other models to improve its performance. Additional visualizations of the qualitative results (D)-(L) are provided in Supplementary Figs. S1-S3.
@@ 4 Discussion
The evaluation of Cellpose, StarDist, and CellViT reveals a persistent performance gap between general and kidney-specific nuclei segmentation, indicating that domain-specific tuning is still required. Our foundation model-based data enrichment strategies aim to meet the need for a kidney-level cell foundation model and to reduce labeling costs. In this process, the collective image curation from multiple foundation models is crucial, as a single model may not capture the diversity of segmentation styles in the target kidney domain. By leveraging the versatility of multiple foundation models, annotation efforts are minimized and transformed into time-efficient image rating curation, which can be further enhanced through the model’s uncertainty analysis. In the first round of HITL fine-tuning, all three foundation models demonstrated improved performance over their baselines, with StarDist achieving the highest segmentation performance, achieving an F1 score of 0.82. The combined use of “easy” (pseudo-labeled) and “hard” (pathologist-corrected) patches proved to be the most effective and broadly applicable strategy, consistently improving performance across all three models. Although Cellpose shows a significant increase in performance (with the F1 score rising from 0.67 to 0.75) with our fine-tuning, it demonstrates lower efficacy when only hard image samples are included. This is primarily because the Cellpose model is primarily designed for fluorescently labeled images (Fluorescent-dominant model). For the training of segmentation task, 512 $\times$ 512 images were converted into grayscale nuclear bands, which can lead to information loss, particularly for hard images with very light staining colors and complex tissue morphological patterns. In contrast, the other two models have been adapted for H&E-stained images.
To clarify the definition of pathology-related foundation models in our study, the three models are image-only foundation models, highlighting scale, data modalities (incorporating different microscopy techniques). Among three models, CellViT is a transformer-based model built upon DINO/SAM and regarded as a foundation model in recent works [44, 45, 30]. Cellpose and StarDist are CNN-based generalist models. CNN-based models that are pre-trained with large-scale representative data are also regarded as foundation models in [46, 47].
While applying multiple foundation models leverages each model’s specialty and reduces annotation costs, our data enrichment and continual learning strategy still heavily relies on pseudo-labels, a weakly supervised approach. Thus, label quality may introduce noise. As shown in Table 2 and Fig. 6, baseline model performance varies. For instance, although both the Cellpose and StarDist models contribute a similar number of “easy” (i.e., “good”) patches, the inferior baseline model (e.g., Cellpose for our kidney images) might produce lower-quality pseudo-labels compared to others. Additionally, since the image-level rating category does not fully capture local performance, it cannot completely reflect mask quality. These factors suggest that disparities can introduce noise into the fine-tuning dataset.
Biases can be introduced by the rating system. (1) Limited rating categories (“good”, “medium”, “bad”). We quantify the percentage of correctly predicted nuclei per category, yet the “medium” class often exhibits the highest variance, as demonstrated in Fig. 5’s cross-model evaluation. This ambiguity can lead to inconsistent ratings and uncertainty. (2) Inter-rater variability may stem from subjective observation practices, particularly since nuclei are small and their boundaries are indistinct. Even with a pathologist’s final review, these biases cannot be entirely eliminated. (3) Image-level ratings may overlook fine-grained details in patch predictions. Therefore, introducing local-level ratings, along with improved rating categories, could enhance the system’s generalizability and robustness.
@@
As the models become better adapted to the specific kidney organ and the performance gap decreases among three models after the first round of fine-tuning, we anticipate generating improved curation results, including a higher number of “good” ratings and more accurate prediction masks in the next round of HITL fine-tuning. Additionally, improved accuracy in nuclei instance segmentation is critical for the extraction and reliability of cell pathomic features (e.g., size, shape, area). As in our previous studies [48, 49], the future work will involve implementing a joint training scheme that combines classification with segmentation, enabling the classification task to guide the segmentation process.
In conclusion, our evaluation of a comprehensive kidney dataset shows that cell nuclei segmentation in histopathology requires further improvement through more organ-targeted foundation models (e.g., kidney-specific). The proposed human-in-the-loop data enrichment strategy—combining predictions from multiple models with limited expert corrections on challenging cases—consistently enhances performance across all models. Improved organ-specific foundation models can further increase the accuracy and reliability of cell pathomic feature extraction (e.g., size, shape, area) and can be integrated into QuPath [50] software to support more efficient workflows in clinical kidney pathology.
5 Data availability
The source data that support the plots within this manuscript is included in the Supplementary Data. The raw data is available upon reasonable request to the corresponding author [Y.H.] (yuankai.huo@vanderbilt.edu) and after satisfying Vanderbilt University and Vanderbilt University Medical Center’s data use agreement.
6 Code availability
@@
Code will be made publicly available at the AFM_kidney_cells repository [51].
7 Supplementary Information
Supplementary materials are provided in the AFM_kidney_cells repository [51] within the “stage2_paper” directory.
@@ Acknowledgments
This work was supported by the National Institutes of Health under award numbers R01EB017230, R01DK135597, T32EB001628, K01AG073584, and 5T32GM007347, DoD HT9425-23-1-0003 (HCY), and in part by the National Center for Research Resources and Grant UL1 RR024975-01. This study was also supported by the National Science Foundation (1452485, 1660816, and 1750213). The Vanderbilt Institute for Clinical and Translational Research (VICTR) is funded by the National Center for Advancing Translational Sciences (NCATS) Clinical Translational Science Award (CTSA) Program, Award Number 5UL1TR002243- 03. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH or NSF. This work was also supported by Vanderbilt Seed Success Grant, Vanderbilt Discovery Grant, and VISE Seed Grant. We extend gratitude to NVIDIA for their support by means of the NVIDIA hardware grant. This work was also supported by NSF NAIRR Pilot Award NAIRR240055.
@@
Author Contributions
Conceptualization: J.G., C.C., H.Y., Y.H.; Study design: J.G., R.D., H.Y., Y.H.; Investigation: J.G., C.C., Y.H.; Data curation and annotation: J.G., S.L., Z.T., Y.L., M.Y., H.Y., Y.H.; Visualization: J.G., S.L., R.D., H.Y., Y.H.; Writing: J.G., S.L., H.Y., Y.H.; Review and approve final manuscript: S.L., C.C., R.D., T.Y., M.L., Q.L., J.X., Y.W., S.Z., C.C. (Catie Chang), M.W., A.F., M.Y., H.Y., Y.H.; Funding acquisition: H.Y., Y.H.; Supervision: Y.H.;
Correspondence to Yuankai Huo (yuankai.huo@vanderbilt.edu)